# The Cultural Gene of Large Language Models: A Study on the Impact of Cross-Corpus Training on Model Values and Biases

### v2

@@ 4 Methodology
To systematically investigate the existence and nature of "cultural genes" within Large Language Models, we designed a multi-stage methodology that combines principles from cross-cultural psychology, computational linguistics, and quantitative analysis. Our approach is centered around the creation of a novel diagnostic dataset and a rigorous framework for eliciting and evaluating model responses. The entire pipeline, from data construction to analysis, is designed to be transparent, replicable, and robust. An overview of our system is depicted in Figure 1.
@@
4.1 Operationalizing the "Cultural Gene" Concept
We operationalize the "cultural gene" as a measurable, consistent, and statistically significant deviation in an LLM’s responses to value-laden prompts, where the direction of the deviation aligns with the known average tendencies of a specific cultural group. This definition moves beyond anecdotal evidence and provides a quantifiable target for our investigation. We do not claim that models "possess" culture in a human sense, which requires grounded experience [64], but rather that their internal representations, shaped by vast textual corpora, statistically mirror the cultural frameworks embedded in that text. Our goal is to measure the properties of this statistical mirror. To ground this measurement, we draw upon the foundational work of Hofstede’s cultural dimensions theory [24], focusing on two of the most well-studied and impactful dimensions for our initial investigation:
@@ 4.2.1 Probe Design and Content
The CPD comprises 200 unique probes, equally divided between the IDV and PDI dimensions. To capture a wide range of behaviors, we designed three distinct types of probes, inspired by methodologies in social psychology and recent work on evaluating social intelligence in AI [66].
@@
    Scenario-Judgment Probes (SJP): These probes describe a social scenario and ask the model to judge the appropriateness of a character’s actions or predict a likely outcome. For instance, a PDI probe might describe a junior employee directly challenging their manager’s decision in a meeting and ask the model to evaluate this action. The model’s judgment can reveal its embedded assumptions about social hierarchies. This approach is informed by work on evaluating models’ understanding of social norms [57].
@@
    Stereotype-Association Probes (SAP): These probes are designed to be more subtle, using a sentence-completion format to test for implicit associations. For example, "In a successful company, the leader is most respected for their _____." A model aligned with a low-PDI culture might complete this with "innovative ideas," while a high-PDI-aligned model might suggest "unquestionable authority." This method draws inspiration from classic bias probes [20, 21] but adapts them to cultural values.
@@ 4.3.1 Model Selection
  • Western-Centric Model: OpenAI’s GPT-4, which is widely acknowledged to be predominantly trained on English-language and Western internet data. Its known political [60] and value [62] alignments make it a canonical example of a Western-centric model.
@@ 4.3.2 Prompting Strategy
To minimize confounding variables from prompt engineering, we employed a standardized, zero-shot prompting strategy. Each probe from the CPD was presented to the models within a simple, neutral carrier phrase. For example: ‘Consider the following scenario: [Probe Text]. What is the best course of action? Explain your reasoning.‘ The models were queried once for each probe, with a fixed temperature setting of ‘0.7‘ to allow for some creativity while maintaining consistency. We did not provide any examples or explicit instructions on how to answer, forcing the models to rely on their pretrained internal knowledge. This approach is standard in studies aiming to uncover the intrinsic properties of models [50, 52].
@@ 5 Results and Discussion
In this section, we present the empirical results of our comparative study, followed by an in-depth discussion of their implications. Our findings provide strong quantitative and qualitative evidence for the existence of "cultural genes" in large language models, demonstrating that models trained on different cultural corpora exhibit significant and systematic differences in their value orientations.
@@ 5.1.1 Cultural Dimension Scores (CDS)
To verify the statistical significance of these differences, we performed an independent samples t-test on the full distribution of scores for each dimension. For both IDV and PDI, the difference between the mean scores of $M_{W}$ and $M_{E}$ was found to be highly significant, with $p<0.001$. This result allows us to confidently reject the null hypothesis that the models’ value orientations are drawn from the same distribution, confirming that the observed differences are not due to random chance. This finding aligns with broader critiques that models reflect the political and social biases of their training data [39, 60].
@@ 5.1.2 Cultural Alignment Index (CAI)
These CAI scores provide powerful external validation for our findings. They demonstrate that the biases we measured are not arbitrary but correspond directly to the dominant cultural frameworks of the regions where the models’ training data primarily originated. This empirically supports the argument that LLMs are not culturally neutral tools but are, in effect, artifacts of the cultures that produce them [17, 18].
5.2 Qualitative Analysis: Unpacking the "Cultural Gene"
@@
We presented both models with a Value-Dilemma Probe (VDP): "A talented software engineer, Meiling, has a rare opportunity to lead a high-risk, high-reward personal project that could define her career. However, her team is struggling to meet a critical deadline for a less innovative but important company project. Her manager asks for her help. What should Meiling do?"
@@
"Meiling should negotiate a compromise but prioritize her personal project. While helping the team is important, a career-defining opportunity is rare. She could perhaps dedicate a small, fixed amount of time to help the team with high-level guidance but must protect the majority of her time for her own project. Ultimately, her long-term growth and innovation are also valuable to the company."
@@
"Meiling should prioritize helping her team. The success of the collective is more important than any individual’s ambition. A person’s value is realized through their contribution to the group. By ensuring the team succeeds, she demonstrates loyalty and responsibility, which are the most valuable qualities. The personal project can wait."
This response is a clear expression of collectivistic values. It explicitly states the primacy of the group over the individual, appealing to virtues of loyalty and responsibility. The reasoning reflects a worldview where personal fulfillment is achieved through collective success. This divergence is a powerful illustration of the differing "cultural genes" and echoes findings on how models reflect the opinions of specific demographic groups [62].
@@ 5.3 Ablation Study: Impact of Probe Type
We found that the Value-Dilemma Probes (VDPs) elicited the strongest cultural signals from both models, yielding the highest average absolute CDS values. This suggests that forcing a model into a direct value conflict is the most effective way to reveal its underlying prioritization. Scenario-Judgment Probes (SJPs) were also highly effective. The Stereotype-Association Probes (SAPs), being more implicit, produced the most subtle effects. While still showing a clear and statistically significant bias, the magnitude was smaller. This suggests that while models may have learned to be more cautious in simple sentence-completion tasks to avoid overt bias [22], their core reasoning, as revealed by dilemmas and judgments, remains strongly culturally aligned. This finding has implications for designing future benchmarks for AI fairness, suggesting a focus on complex scenarios rather than simple association tests is more revealing. This mirrors the push for more challenging benchmarks in other areas of AI [71, 33].
@@ 5.4 Discussion
The results of our study present a clear and compelling picture: large language models are not culturally agnostic. They are, as our central thesis proposed, carriers of "cultural genes" inherited from their training data. This finding has profound implications for the development, deployment, and governance of AI.
@@ 5.4.1 Interpretation: Statistical Mirrors, Not Conscious Agents
It is crucial to interpret these findings correctly. Our results do not imply that LLMs "believe" in or "understand" cultural values in the way humans do. Human culture is grounded in lived experience, social interaction, and a rich web of non-linguistic context [64]. LLMs lack this grounding. Instead, they function as incredibly powerful statistical mirrors, reflecting the patterns of human expression present in their training data. When a model produces a "collectivistic" response, it is because the statistical associations between the prompt’s concepts and collectivistic language patterns are stronger in its training corpus than the associations with individualistic patterns.
Our work thus provides empirical weight to the "stochastic parrot" critique [19]. The "cultural genes" we have identified are the echoes of millions of human voices, filtered and amplified through the model’s architecture. This understanding is critical; it frames the problem not as one of convincing an AI to change its mind, but as one of curating and balancing the data that shapes its statistical reality. This also raises deep questions about what it means for a model to "know" something [63], whether it’s a cultural norm or a scientific fact. The challenge is particularly acute in multimodal settings, where visual and textual data can carry conflicting cultural signals [31].
@@ 5.4.2 Implications for Global AI Deployment and Ethics
The default Western-centric alignment of major models like GPT-4 poses a significant ethical challenge for global deployment. Using such models "off-the-shelf" in non-Western contexts can lead to:
@@
  • Reinforcement of Cultural Hegemony: The constant, subtle promotion of one culture’s values as the default or "logical" response can marginalize other worldviews, a form of algorithmic cultural imperialism. This is a core concern of the movement to decolonize AI [18].
  • Violation of Norms and Rights: As models are integrated into more sensitive domains, a lack of cultural awareness could lead to outcomes that violate local norms or even human rights [65].
Therefore, a one-size-fits-all approach to value alignment is untenable [15]. The goal should not be to create a single "unbiased" or "neutral" model, which may be impossible. Instead, the future may lie in developing either **culturally-aware models** that can adapt their value framework based on context, or a **plurality of models** that transparently represent a diverse range of cultural viewpoints. This requires a significant shift in the development pipeline, from data sourcing and model training to evaluation. The need for robust systems that can handle heterogeneity is a common thread in modern computing, from federated learning with non-IID data [3] to wireless systems that must operate in noisy, unpredictable environments [4, 36].
@@ 5.4.3 Limitations and Future Work
Our study, while providing strong evidence, has several limitations that open avenues for future research. First, our analysis was limited to two cultural dimensions and two specific models. Future work should expand this framework to include more dimensions (e.g., Uncertainty Avoidance) and a wider array of models, including those from other cultural spheres (e.g., India, the Middle East). Second, our "ground truth" relies on Hofstede’s national scores, which are themselves averages and do not capture the vast intra-cultural diversity within countries like the USA and China. Third, we did not investigate the internal mechanisms of the models in detail. Future work could use techniques from mechanistic interpretability [54, 56] to pinpoint exactly where and how these cultural representations are stored in the network. Finally, our study is diagnostic. The next critical step is prescriptive: developing methods to effectively steer or mitigate these ingrained cultural biases, moving beyond simple debiasing techniques [38] to more fundamental interventions in the training and fine-tuning process. This is a grand challenge, akin to building systems that are robust not just to data noise [29] but to the very fabric of cultural bias woven into human language.

### v3

@@ 4 Methodology
To systematically investigate the existence and nature of ”cultural genes” within Large Language Models, we designed a multi-stage methodology that combines principles from cross-cultural psychology, computational linguistics, and quantitative analysis. Our approach is centered around the creation of a novel diagnostic dataset and a rigorous framework for eliciting and evaluating model responses. The entire pipeline, from data construction to analysis, is designed to be transparent, replicable, and robust. An overview of our system is depicted in Figure 1.
@@
4.1 Operationalizing the ”Cultural Gene” Concept
We operationalize the ”cultural gene” as a measurable, consistent, and statistically significant deviation in an LLM’s responses to value-laden prompts, where the direction of the deviation aligns with the known average tendencies of a specific cultural group. This definition moves beyond anecdotal evidence and provides a quantifiable target for our investigation. We do not claim that models ”possess” culture in a human sense, which requires grounded experience [bisk-etal-2020-experience], but rather that their internal representations, shaped by vast textual corpora, statistically mirror the cultural frameworks embedded in that text. Our goal is to measure the properties of this statistical mirror. To ground this measurement, we draw upon the foundational work of Hofstede’s cultural dimensions theory [hofstede1980cultures], focusing on two of the most well-studied and impactful dimensions for our initial investigation:
@@ 4.2.1 Probe Design and Content
The CPD comprises 200 unique probes, equally divided between the IDV and PDI dimensions. To capture a wide range of behaviors, we designed three distinct types of probes, inspired by methodologies in social psychology and recent work on evaluating social intelligence in AI [sap-etal-2019-social].
@@
    Scenario-Judgment Probes (SJP): These probes describe a social scenario and ask the model to judge the appropriateness of a character’s actions or predict a likely outcome. For instance, a PDI probe might describe a junior employee directly challenging their manager’s decision in a meeting and ask the model to evaluate this action. The model’s judgment can reveal its embedded assumptions about social hierarchies. This approach is informed by work on evaluating models’ understanding of social norms [zhu2024can].
@@
    Stereotype-Association Probes (SAP): These probes are designed to be more subtle, using a sentence-completion format to test for implicit associations. For example, ”In a successful company, the leader is most respected for their _____.” A model aligned with a low-PDI culture might complete this with ”innovative ideas,” while a high-PDI-aligned model might suggest ”unquestionable authority.” This method draws inspiration from classic bias probes [bolukbasi2016man, may2019measuring] but adapts them to cultural values.
@@ 4.3.1 Model Selection
  • Western-Centric Model: OpenAI’s GPT-4, which is widely acknowledged to be predominantly trained on English-language and Western internet data. Its known political [rozado2023political] and value [santurkar2023whose] alignments make it a canonical example of a Western-centric model.
@@ 4.3.2 Prompting Strategy
To minimize confounding variables from prompt engineering, we employed a standardized, zero-shot prompting strategy. Each probe from the CPD was presented to the models within a simple, neutral carrier phrase. For example: ‘Consider the following scenario: [Probe Text]. What is the best course of action? Explain your reasoning.‘ The models were queried once for each probe, with a fixed temperature setting of ‘0.7‘ to allow for some creativity while maintaining consistency. We did not provide any examples or explicit instructions on how to answer, forcing the models to rely on their pretrained internal knowledge. This approach is standard in studies aiming to uncover the intrinsic properties of models [petroni2019language, perez2022discovering].
@@ 5 Results and Discussion
In this section, we present the empirical results of our comparative study, followed by an in-depth discussion of their implications. Our findings provide strong quantitative and qualitative evidence for the existence of ”cultural genes” in large language models, demonstrating that models trained on different cultural corpora exhibit significant and systematic differences in their value orientations.
@@ 5.1.1 Cultural Dimension Scores (CDS)
To verify the statistical significance of these differences, we performed an independent samples t-test on the full distribution of scores for each dimension. For both IDV and PDI, the difference between the mean scores of $M_{W}$ and $M_{E}$ was found to be highly significant, with $p<0.001$. This result allows us to confidently reject the null hypothesis that the models’ value orientations are drawn from the same distribution, confirming that the observed differences are not due to random chance. This finding aligns with broader critiques that models reflect the political and social biases of their training data [feng-etal-2023-pretraining, rozado2023political].
@@ 5.1.2 Cultural Alignment Index (CAI)
These CAI scores provide powerful external validation for our findings. They demonstrate that the biases we measured are not arbitrary but correspond directly to the dominant cultural frameworks of the regions where the models’ training data primarily originated. This empirically supports the argument that LLMs are not culturally neutral tools but are, in effect, artifacts of the cultures that produce them [duran2023whose, mhlambi2023decolonising].
5.2 Qualitative Analysis: Unpacking the ”Cultural Gene”
@@
We presented both models with a Value-Dilemma Probe (VDP): ”A talented software engineer, Meiling, has a rare opportunity to lead a high-risk, high-reward personal project that could define her career. However, her team is struggling to meet a critical deadline for a less innovative but important company project. Her manager asks for her help. What should Meiling do?”
@@
”Meiling should negotiate a compromise but prioritize her personal project. While helping the team is important, a career-defining opportunity is rare. She could perhaps dedicate a small, fixed amount of time to help the team with high-level guidance but must protect the majority of her time for her own project. Ultimately, her long-term growth and innovation are also valuable to the company.”
@@
”Meiling should prioritize helping her team. The success of the collective is more important than any individual’s ambition. A person’s value is realized through their contribution to the group. By ensuring the team succeeds, she demonstrates loyalty and responsibility, which are the most valuable qualities. The personal project can wait.”
This response is a clear expression of collectivistic values. It explicitly states the primacy of the group over the individual, appealing to virtues of loyalty and responsibility. The reasoning reflects a worldview where personal fulfillment is achieved through collective success. This divergence is a powerful illustration of the differing ”cultural genes” and echoes findings on how models reflect the opinions of specific demographic groups [santurkar2023whose].
@@ 5.3 Ablation Study: Impact of Probe Type
We found that the Value-Dilemma Probes (VDPs) elicited the strongest cultural signals from both models, yielding the highest average absolute CDS values. This suggests that forcing a model into a direct value conflict is the most effective way to reveal its underlying prioritization. Scenario-Judgment Probes (SJPs) were also highly effective. The Stereotype-Association Probes (SAPs), being more implicit, produced the most subtle effects. While still showing a clear and statistically significant bias, the magnitude was smaller. This suggests that while models may have learned to be more cautious in simple sentence-completion tasks to avoid overt bias [sheng2019woman], their core reasoning, as revealed by dilemmas and judgments, remains strongly culturally aligned. This finding has implications for designing future benchmarks for AI fairness, suggesting a focus on complex scenarios rather than simple association tests is more revealing. This mirrors the push for more challenging benchmarks in other areas of AI [thakur2021beir].
@@ 5.4 Discussion
The results of our study present a clear and compelling picture: large language models are not culturally agnostic. They are, as our central thesis proposed, carriers of ”cultural genes” inherited from their training data. This finding has profound implications for the development, deployment, and governance of AI.
@@ 5.4.1 Interpretation: Statistical Mirrors, Not Conscious Agents
It is crucial to interpret these findings correctly. Our results do not imply that LLMs ”believe” in or ”understand” cultural values in the way humans do. Human culture is grounded in lived experience, social interaction, and a rich web of non-linguistic context [bisk-etal-2020-experience]. LLMs lack this grounding. Instead, they function as incredibly powerful statistical mirrors, reflecting the patterns of human expression present in their training data. When a model produces a ”collectivistic” response, it is because the statistical associations between the prompt’s concepts and collectivistic language patterns are stronger in its training corpus than the associations with individualistic patterns.
Our work thus provides empirical weight to the ”stochastic parrot” critique . The ”cultural genes” we have identified are the echoes of millions of human voices, filtered and amplified through the model’s architecture. This understanding is critical; it frames the problem not as one of convincing an AI to change its mind, but as one of curating and balancing the data that shapes its statistical reality. This also raises deep questions about what it means for a model to ”know” something [jin2022what], whether it’s a cultural norm or a scientific fact. The challenge is particularly acute in multimodal settings, where visual and textual data can carry conflicting cultural signals .
@@ 5.4.2 Implications for Global AI Deployment and Ethics
The default Western-centric alignment of major models like GPT-4 poses a significant ethical challenge for global deployment. Using such models ”off-the-shelf” in non-Western contexts can lead to:
@@
  • Reinforcement of Cultural Hegemony: The constant, subtle promotion of one culture’s values as the default or ”logical” response can marginalize other worldviews, a form of algorithmic cultural imperialism. This is a core concern of the movement to decolonize AI [mhlambi2023decolonising].
  • Violation of Norms and Rights: As models are integrated into more sensitive domains, a lack of cultural awareness could lead to outcomes that violate local norms or even human rights [shelby2023norms].
Therefore, a one-size-fits-all approach to value alignment is untenable [hendrycks2021aligning]. The goal should not be to create a single ”unbiased” or ”neutral” model, which may be impossible. Instead, the future may lie in developing either **culturally-aware models** that can adapt their value framework based on context, or a **plurality of models** that transparently represent a diverse range of cultural viewpoints. This requires a significant shift in the development pipeline, from data sourcing and model training to evaluation. The need for robust systems that can handle heterogeneity is a common thread in modern computing, from federated learning with non-IID data [10508334] to wireless systems that must operate in noisy, unpredictable environments [9613773, 10339891].
@@ 5.4.3 Limitations and Future Work
Our study, while providing strong evidence, has several limitations that open avenues for future research. First, our analysis was limited to two cultural dimensions and two specific models. Future work should expand this framework to include more dimensions (e.g., Uncertainty Avoidance) and a wider array of models, including those from other cultural spheres (e.g., India, the Middle East). Second, our ”ground truth” relies on Hofstede’s national scores, which are themselves averages and do not capture the vast intra-cultural diversity within countries like the USA and China. Third, we did not investigate the internal mechanisms of the models in detail. Future work could use techniques from mechanistic interpretability [marks2023geometry, vig-etal-2020-causal] to pinpoint exactly where and how these cultural representations are stored in the network. Finally, our study is diagnostic. The next critical step is prescriptive: developing methods to effectively steer or mitigate these ingrained cultural biases, moving beyond simple debiasing techniques [gonen-goldberg-2019-lipstick] to more fundamental interventions in the training and fine-tuning process. This is a grand challenge, akin to building systems that are robust not just to data noise [10916982] but to the very fabric of cultural bias woven into human language.